OpenStudy (anonymous):

The length of a rectangle is twice its width. If the area of the rectangle is 200 yd 2, find its perimeter.

OpenStudy (anonymous):

60 yds =2*20+2*10

OpenStudy (shinalcantara):

Area for a rectangle is given by the formula: A=l*w You're given the area, A=200 sq yds And that length, l=2w Substituting it to the original formula you'll have: A = l*w A = 2w (w) A = 2w^2 200 = 2 w^2 w^2 = 100 w = 100 yrds Since width = 100 yrds then length = 2w l = 2(100) l = 200 yards Perimeter = 2L + 2W = 2(200) + 2(100) = 600 yards